# PoLAR: Polar-Decomposed Low-Rank Adapter Representation

### v1

@@ 5 Conclusion
Appendix A Future Directions, Limitations, and Broader Impact
Future Direction: Spectral Distribution in Low-Rank Learning. Given the low stable rank of the LoRA update, a closer examination of the spectral dynamics of low-rank learning seems to be a promising direction for future work. While several works indicate that full-rank neural network training is already biased towards learning low-rank changes to the initial weight (Boix-Adsera et al., 2023; Mousavi-Hosseini et al., 2023), the impact of the explicit low-rank constraint on these dynamics is not yet well understood. Potentially, the low-rank constraint could act similar to the scenario of over-regularization in Martin and Mahoney (2021), leading to a significant rank collapse. Similarly, the low-rank constraint could alter the incremental learning dynamics whereby the difference between the trained and initial weights progressively increases in stable rank (Boix-Adsera et al., 2023). An investigation into the dynamics of the spectral distributions of PoLAR could bring more insights into the benefits of the parameterization.
@@ Appendix B Additional Related Work
Parameter-Efficient Fine-Tuning. The goal of PEFT is to reduce the resource requirement for finetuning LLMs on downstream tasks. Some commonly employed methods include adapters (Houlsby et al., 2019) and prefix tuning (Li and Liang, 2021). The family of LoRA adapters (Hu et al., 2022) recently gained significant popularity. Beyond the works already mentioned in Section 1.1, there are several lines of work focused on improving or understanding different aspects of LoRA. Zhu et al. (2024); Hayou et al. (2024a) contribute to the understanding of the different roles of the low-rank factors. Gu et al. (2023) extend the framework to diffusion models. Xu et al. (2024a); Dettmers et al. (2023) improve memory-efficiency through quantization strategies. Liu et al. (2022) further improve parameter efficiency. Zhang et al. (2024) apply landing methods to LoRA fine-tuning and present LoRA variants that constrain the up-projection to the Oblique and Stiefel manifold. Scale and rotation invariance is considered for efficiently optimizing vanilla LoRA parameterization (Yen et al., 2025; Zhang et al., 2025).
Optimization for Matrix Factorization. Existing works mostly handle matrix factorization problems under the BM parameterization, i.e., (3). Such problems are classical examples for understanding the optimization dynamics of LoRA. They involve a complex landscape characterized by nonconvexity, saddle points, and the absence of the PL condition. Recent works have examined the convergence of several algorithms under the BM parameterization, such as GD, Alternating GD, and ScaledGD (Du et al., 2018; Ward and Kolda, 2023; Li et al., 2025) in overparameterized settings. Another closely related problem is matrix sensing. The convergence behaviors of different algorithms under the BM parameterization can be found in e.g., (Stöger and Soltanolkotabi, 2021; Zhang et al., 2021; Jin et al., 2023; Xiong et al., 2024).
@@ Appendix C Useful Facts
C.1 Polar Decomposition
We provide the definition for the polar decomposition below and refer the interested reader to Section 9.4.3 of Golub and Van Loan (2013) for further details.
Definition 1 (Polar Decomposition).
The (right) polar decomposition of a matrix $\mathbf{X}\in\mathbb{R}^{m\times r}$ with $m\geq r$ is defined as $\mathbf{X}=\mathbf{U}\mathbf{P}$ where $\mathbf{U}\in\mathbb{R}^{m\times r}$ has orthonormal columns and $\mathbf{P}\in\mathbb{R}^{r\times r}$ is positive semi-definite.
When $r=1$, the polar decomposition in Definition 1 reduces to the familiar magnitude-direction decomposition of vectors. In general, the polar decomposition can be viewed as an extension to matrices, where $\mathbf{U}$ represents the directional component and $\mathbf{P}$ captures the magnitude.
C.2 Angles Between Subspaces
@@
C.3 Other Useful Lemmas
@@
=$\stackrel{{\scriptstyle(a)}}{{=}}\min_{\|\mathbf{x}\|=1,\|\mathbf{y}\|=1,\mathbf{x}\in\text{ColSpan}(\mathbf{B}),\mathbf{y}\in\text{ColSpan}(\mathbf{B})}\|\mathbf{A}\mathbf{y}\|\cdot\|\mathbf{B}\mathbf{x}\|$
@@
C.4 Equivalence of Matrix Factorization and LoRA
Whitening data refers to transforming the data such that the empirical uncentered covariance matrix of the features is identity. Suppose $\mathbf{D}\in\mathbb{R}^{n\times N}$ holds $N$ training examples in its columns with $n$ features each, then whitened data refers to having $\mathbf{D}\mathbf{D}^{\top}=\mathbf{I}_{n}$. We follow standard arguments laid out in Arora et al. (2018); Li et al. (2025) to show that low-rank adaptation on a linear model with whitened data and squared loss can be written as a matrix factorization problem. Consider the minimization of
@@ C.4 Equivalence of Matrix Factorization and LoRA
=$\overset{(a)}{=}\mathsf{Tr}((\mathbf{W}_{0}+\mathbf{X}\mathbf{Y}^{\top})(\mathbf{W}_{0}+\mathbf{X}\mathbf{Y}^{\top})^{\top})-\mathsf{Tr}((\mathbf{W}_{0}+\mathbf{X}\mathbf{Y}^{\top})\mathbf{D}\mathbf{A}^{\top})$
@@
=$\overset{(b)}{=}\mathsf{Tr}(((\mathbf{W}_{0}+\mathbf{X}\mathbf{Y}^{\top})-\bm{\Lambda})((\mathbf{W}_{0}+\mathbf{X}\mathbf{Y}^{\top})-\bm{\Lambda})^{\top})-\mathsf{Tr}(\bm{\Lambda}\bm{\Lambda}^{\top})+\mathsf{Tr}(\mathbf{A}\mathbf{A}^{\top})$
@@ D.2 General Dynamics
=$\stackrel{{\scriptstyle(a)}}{{=}}\big{[}\mathbf{I}_{r_{A}}+\eta(\mathbf{I}_{r_{A}}-\mathbf{\Phi}_{t}\mathbf{\Phi}_{t}^{\top})\mathbf{\Sigma}\mathbf{\Psi}_{t}\mathbf{\Psi}_{t}^{\top}\mathbf{\Sigma}\big{]}\mathbf{\Phi}_{t}\big{(}\mathbf{I}_{r}+\eta^{2}\mathbf{E}_{t}^{\top}\mathbf{E}_{t}\big{)}^{-1/2}$
@@
=$\stackrel{{\scriptstyle(b)}}{{=}}\big{[}\mathbf{I}_{r_{A}}+\eta(\mathbf{I}_{r_{A}}-\mathbf{\Psi}_{t}\mathbf{\Psi}_{t}^{\top})\mathbf{\Sigma}\mathbf{\Phi}_{t}\mathbf{\Phi}_{t}^{\top}\mathbf{\Sigma}\big{]}\mathbf{\Psi}_{t}\big{(}\mathbf{I}_{r}+\eta^{2}\mathbf{F}_{t}^{\top}\mathbf{F}_{t}\big{)}^{-1/2}$
@@ D.5.1 Non-Increasing Misalignment of $\mathbf{X}_{t}$
=$\stackrel{{\scriptstyle(a)}}{{=}}\big{(}\mathbf{\Omega}_{t}-\eta\mathbf{U}_{\perp}^{\top}\mathbf{X}_{t}\mathbf{\Theta}_{t}\mathbf{\Theta}_{t}^{\top}\big{)}(\mathbf{I}_{r}+\eta^{2}\mathbf{E}_{t}^{\top}\mathbf{E}_{t})^{-1/2}$
@@ D.6.1 Dynamics of $\mathsf{Tr}(\mathbf{\Phi}_{t}\mathbf{\Phi}_{t}^{\top})$
Phase II. $0.50.5<\mathsf{Tr}(\mathbf{I}_{r_{A}}-\mathbf{\Phi}_{t}\mathbf{\Phi}_{t}^{\top})<{r_{A}}-$ 0.5. Suppose that $\mathsf{Tr}\big{(}(\mathbf{I}_{r_{A}}-\mathbf{\Lambda}_{t})\mathbf{\Lambda}_{t}\big{)}\geq\rho$, for some $\rho>0$ to be discussed shortly. Let $\eta={\cal O}\big{(}\frac{\rho(r-r_{A})^{2}}{r^{2}\kappa^{2}m}\big{)}$ and $\eta\leq 0.5$, it is straightforward to have
@@
Next, we show that $\rho\geq{\cal O}(\min\{\frac{1}{m},\frac{(r-r_{A})^{2}}{mr}\})$. Notice that $\mathsf{Tr}\big{(}(\mathbf{I}_{r_{A}}-\mathbf{\Lambda}_{t})\mathbf{\Lambda}_{t}\big{)}\geq\sum_{i=1}^{r_{A}}\sigma_{i}^{2}(\mathbf{\Phi}_{t})(1-\sigma_{i}^{2}(\mathbf{\Phi}_{t}))\geq\sigma_{r_{A}}^{2}(\mathbf{\Phi}_{t})(1-\sigma_{r_{A}}^{2}(\mathbf{\Phi}_{t}))\geq{\cal O}(\min\{\frac{1}{m},\frac{(r-r_{A})^{2}}{mr}\})$, where the last inequality comes from the facts that i) for $x\in[a,b]$ with $10<a<0.5<b<$ 1, the smallest value of $x(1-x)$ is $\min\{a(1-a),b(1-b)\}$; and, ii) $\sigma_{r_{A}}^{2}(\mathbf{\Phi}_{t})$ belongs to interval $[a,b]$ with $a={\cal O}\big{(}\frac{(r-r_{A})^{2}}{mr})$ and $b=\frac{r_{A}-0.5}{r_{A}}=1-\frac{1}{2r_{A}}\leq 1-\frac{1}{m}$. Lemmas 14 and 11 are adopted to calculate $a$, that is, $\sigma_{r_{A}}^{2}(\mathbf{\Phi}_{t})\geq\sigma_{r_{A}}^{2}(\mathbf{\Phi}_{0})={\cal O}((r-r_{A})^{2}/mr)$.
@@ D.6.2 Dynamics of $\mathsf{Tr}(\mathbf{\Psi}_{t}\mathbf{\Psi}_{t}^{\top})$
Phase II. $0.50.5<\mathsf{Tr}(\mathbf{I}_{r_{A}}-\mathbf{\Psi}_{t}\mathbf{\Psi}_{t}^{\top})<{r_{A}}-$ 0.5. Suppose that $\mathsf{Tr}\big{(}(\mathbf{I}_{r_{A}}-\mathbf{\Lambda}_{t})\mathbf{\Lambda}_{t}\big{)}\geq\rho$, for some $\rho$ to be discussed shortly. Choosing $\eta\leq 0.5$, and $\eta={\cal O}\big{(}\frac{\rho(r-r_{A})^{2}}{r^{2}\kappa^{2}m}\big{)}$, it is straightforward to have
@@
Next, we show that $\rho\geq{\cal O}(\min\{\frac{1}{n},\frac{(r-r_{A})^{2}}{nr}\})\geq{\cal O}(\min\{\frac{1}{m},\frac{(r-r_{A})^{2}}{mr}\})$. Notice that $\mathsf{Tr}\big{(}(\mathbf{I}_{r_{A}}-\tilde{\mathbf{\Lambda}}_{t})\tilde{\mathbf{\Lambda}}_{t}\big{)}\geq\sum_{i=1}^{r_{A}}\sigma_{i}^{2}(\mathbf{\Psi}_{t})(1-\sigma_{i}^{2}(\mathbf{\Psi}_{t}))\geq\sigma_{r_{A}}^{2}(\mathbf{\Psi}_{t})(1-\sigma_{r_{A}}^{2}(\mathbf{\Psi}_{t}))\geq{\cal O}(\min\{\frac{1}{n},\frac{(r-r_{A})^{2}}{nr}\})$, where the last inequality comes from the facts that i) for $x\in[a,b]$ with $10<a<0.5<b<$ 1, the smallest value of $x(1-x)$ is $\min\{a(1-a),b(1-b)\}$; and, ii) $\sigma_{r_{A}}^{2}(\mathbf{\Psi}_{t})$ belongs to interval $[a,b]$ with $a={\cal O}(\frac{(r-r_{A})^{2}}{nr})$ and $b=\frac{r_{A}-0.5}{r_{A}}=1-\frac{1}{2r_{A}}\leq 1-\frac{1}{n}$. Lemmas 15 and 11 are adopted to calculate $a$, that is, $a=\sigma_{r_{A}}^{2}(\mathbf{\Psi}_{t})\geq\sigma_{r_{A}}^{2}(\mathbf{\Psi}_{0})={\cal O}((r-r_{A})^{2}/nr)$.
@@ E.2.5 Convergence
Phase II. $0.50.5<\mathsf{Tr}(\mathbf{I}_{r_{B}}-\mathbf{\Phi}_{t}\mathbf{\Phi}_{t}^{\top})<{r_{B}}-$ 0.5. Suppose that $\mathsf{Tr}\big{(}(\mathbf{I}_{r_{B}}-\mathbf{\Lambda}_{t}^{2})\mathbf{\Lambda}_{t}^{2}\big{)}\geq\rho$, for some $\rho$ to be discussed shortly. Choosing $\eta\leq 0.5$, and $\eta={\cal O}\big{(}\frac{\rho(r-r_{B})^{2}}{r^{2}\kappa^{2}m}\big{)}$, it is straightforward to have
@@
Next, we show that $\rho\geq{\cal O}(\min\{\frac{1}{m},\frac{(r-r_{B})^{2}}{mr}\})$. Notice that $\mathsf{Tr}\big{(}(\mathbf{I}_{r_{B}}-\mathbf{\Lambda}_{t}^{2})\mathbf{\Lambda}_{t}^{2}\big{)}\geq\sum_{i=1}^{r_{B}}\sigma_{i}^{2}(\mathbf{\Phi}_{t})(1-\sigma_{i}^{2}(\mathbf{\Phi}_{t}))\geq\sigma_{r_{B}}^{2}(\mathbf{\Phi}_{t})(1-\sigma_{r_{B}}^{2}(\mathbf{\Phi}_{t}))\geq{\cal O}(\min\{\frac{1}{m},\frac{(r-r_{B})^{2}}{mr}\})$, where the last inequality comes from the facts that i) for $x\in[a,b]$ with $10<a<0.5<b<$ 1, the smallest value of $x(1-x)$ is $\min\{a(1-a),b(1-b)\}$; and, ii) $\sigma_{r_{B}}^{2}(\mathbf{\Phi}_{t})$ belongs to interval $[a,b]$ with $a={\cal O}\big{(}\frac{(r-r_{B})^{2}}{mr})$ and $b=\frac{r_{B}-0.5}{r_{B}}=1-\frac{1}{2r_{B}}\leq 1-\frac{1}{m}$. Lemmas 21 and 19 are adopted to calculate $a$, that is, $\sigma_{r_{B}}^{2}(\mathbf{\Phi}_{t})\geq\sigma_{r_{B}}^{2}(\mathbf{\Phi}_{0})={\cal O}((r-r_{B})^{2}/mr)$.

### v2

@@ Acknowledgements
Contents
Appendix A Limitations, Broader Impact, and Practical Concerns
@@ Appendix B Additional Related Work
Parameter-Efficient Fine-Tuning. The goal of PEFT is to reduce the resource requirement for finetuning LLMs on downstream tasks. Some commonly employed methods include adapters (Houlsby et al., 2019) and prefix tuning (Li and Liang, 2021). The family of LoRA adapters (Hu et al., 2022) recently gained significant popularity. Beyond the works already mentioned in Section 1.1, there are several lines of work focused on improving or understanding different aspects of LoRA. Zhu et al. (2024); Hayou et al. (2024a) contribute to the understanding of the different roles of the low-rank factors. Gu et al. (2023) extend the framework to diffusion models. Xu et al. (2024a); Dettmers et al. (2023) improve memory-efficiency through quantization strategies. Liu et al. (2022) further improve parameter efficiency. Zhang et al. (2024) apply landing methods to LoRA fine-tuning and present LoRA variants that constrain the up-projection to the Oblique and Stiefel manifold.
Optimization for Matrix Factorization. Existing works mostly handle matrix factorization problems under the BM parameterization, i.e., (3). Such problems are classical examples for understanding the optimization dynamics of LoRA. They involve a complex landscape characterized by nonconvexity, saddle points, and the absence of the PL condition.
Recent works have examined the convergence of several algorithms under the BM parameterization, such as GD, Alternating GD, and ScaledGD (Du et al., 2018; Ward and Kolda, 2023; Li et al., 2025a) in overparameterized settings. Another closely related problem is matrix sensing. The convergence behaviors of different algorithms under the BM parameterization can be found in e.g., (Stöger and Soltanolkotabi, 2021; Zhang et al., 2021; Jin et al., 2023; Xiong et al., 2024).
@@ Appendix C Useful Facts
C.1 Angles Between Subspaces
@@
C.2 Other Useful Lemmas
@@
$\stackrel{{\scriptstyle(a)}}{{=}}\min_{\|\mathbf{x}\|=1,\|\mathbf{y}\|=1,\mathbf{x}\in\text{ColSpan}(\mathbf{B}),\mathbf{y}\in\text{ColSpan}(\mathbf{B})}\|\mathbf{A}\mathbf{y}\|\cdot\|\mathbf{B}\mathbf{x}\|$
@@
C.3 Equivalence of Matrix Factorization and LoRA
Whitening data refers to transforming the data such that the empirical uncentered covariance matrix of the features is identity. Suppose $\mathbf{D}\in\mathbb{R}^{n\times N}$ holds $N$ training examples in its columns with $n$ features each, then whitened data refers to having $\mathbf{D}\mathbf{D}^{\top}=\mathbf{I}_{n}$. We follow standard arguments laid out in Arora et al. (2018); Li et al. (2025a) to show that low-rank adaptation on a linear model with whitened data and squared loss can be written as a matrix factorization problem. Consider the minimization of
@@ C.3 Equivalence of Matrix Factorization and LoRA
$\overset{(a)}{=}\mathsf{Tr}((\mathbf{W}_{0}+\mathbf{X}\mathbf{Y}^{\top})(\mathbf{W}_{0}+\mathbf{X}\mathbf{Y}^{\top})^{\top})-\mathsf{Tr}((\mathbf{W}_{0}+\mathbf{X}\mathbf{Y}^{\top})\mathbf{D}\mathbf{A}^{\top})$
@@
$\overset{(b)}{=}\mathsf{Tr}(((\mathbf{W}_{0}+\mathbf{X}\mathbf{Y}^{\top})-\bm{\Lambda})((\mathbf{W}_{0}+\mathbf{X}\mathbf{Y}^{\top})-\bm{\Lambda})^{\top})-\mathsf{Tr}(\bm{\Lambda}\bm{\Lambda}^{\top})+\mathsf{Tr}(\mathbf{A}\mathbf{A}^{\top})$
@@ D.2 General Dynamics
$\stackrel{{\scriptstyle(a)}}{{=}}\big[\mathbf{I}_{r_{A}}+\eta(\mathbf{I}_{r_{A}}-\mathbf{\Phi}_{t}\mathbf{\Phi}_{t}^{\top})\mathbf{\Sigma}\mathbf{\Psi}_{t}\mathbf{\Psi}_{t}^{\top}\mathbf{\Sigma}\big]\mathbf{\Phi}_{t}\big(\mathbf{I}_{r}+\eta^{2}\mathbf{E}_{t}^{\top}\mathbf{E}_{t}\big)^{-1/2}$
@@
$\stackrel{{\scriptstyle(b)}}{{=}}\big[\mathbf{I}_{r_{A}}+\eta(\mathbf{I}_{r_{A}}-\mathbf{\Psi}_{t}\mathbf{\Psi}_{t}^{\top})\mathbf{\Sigma}\mathbf{\Phi}_{t}\mathbf{\Phi}_{t}^{\top}\mathbf{\Sigma}\big]\mathbf{\Psi}_{t}\big(\mathbf{I}_{r}+\eta^{2}\mathbf{F}_{t}^{\top}\mathbf{F}_{t}\big)^{-1/2}$
@@ D.5.1 Non-Increasing Misalignment of $\mathbf{X}_{t}$
$\stackrel{{\scriptstyle(a)}}{{=}}\big(\mathbf{\Omega}_{t}-\eta\mathbf{U}_{\perp}^{\top}\mathbf{X}_{t}\mathbf{\Theta}_{t}\mathbf{\Theta}_{t}^{\top}\big)(\mathbf{I}_{r}+\eta^{2}\mathbf{E}_{t}^{\top}\mathbf{E}_{t})^{-1/2}$
@@ D.6.1 Dynamics of $\mathsf{Tr}(\mathbf{\Phi}_{t}\mathbf{\Phi}_{t}^{\top})$
Phase II. $0.5<\mathsf{Tr}(\mathbf{I}_{r_{A}}-\mathbf{\Phi}_{t}\mathbf{\Phi}_{t}^{\top})<{r_{A}}-0.5$. Suppose that $\mathsf{Tr}\big((\mathbf{I}_{r_{A}}-\mathbf{\Lambda}_{t})\mathbf{\Lambda}_{t}\big)\geq\rho$, for some $\rho>0$ to be discussed shortly. Let $\eta={\cal O}\big(\frac{\rho(r-r_{A})^{2}}{r^{2}\kappa^{2}m}\big)$ and $\eta\leq 0.5$, it is straightforward to have
@@
Next, we show that $\rho\geq{\cal O}(\min\{\frac{1}{m},\frac{(r-r_{A})^{2}}{mr}\})$. Notice that $\mathsf{Tr}\big((\mathbf{I}_{r_{A}}-\mathbf{\Lambda}_{t})\mathbf{\Lambda}_{t}\big)\geq\sum_{i=1}^{r_{A}}\sigma_{i}^{2}(\mathbf{\Phi}_{t})(1-\sigma_{i}^{2}(\mathbf{\Phi}_{t}))\geq\sigma_{r_{A}}^{2}(\mathbf{\Phi}_{t})(1-\sigma_{r_{A}}^{2}(\mathbf{\Phi}_{t}))\geq{\cal O}(\min\{\frac{1}{m},\frac{(r-r_{A})^{2}}{mr}\})$, where the last inequality comes from the facts that i) for $x\in[a,b]$ with $0<a<0.5<b<1$, the smallest value of x(1x)x(1-x) is $\min\{a(1-a),b(1-b)\}$; and, ii) $\sigma_{r_{A}}^{2}(\mathbf{\Phi}_{t})$ belongs to interval $[a,b]$ with $a={\cal O}\big(\frac{(r-r_{A})^{2}}{mr})$ and $b=\frac{r_{A}-0.5}{r_{A}}=1-\frac{1}{2r_{A}}\leq 1-\frac{1}{m}$. Lemmas 14 and 11 are adopted to calculate $a$, that is, $\sigma_{r_{A}}^{2}(\mathbf{\Phi}_{t})\geq\sigma_{r_{A}}^{2}(\mathbf{\Phi}_{0})={\cal O}((r-r_{A})^{2}/mr)$.
@@ D.6.2 Dynamics of $\mathsf{Tr}(\mathbf{\Psi}_{t}\mathbf{\Psi}_{t}^{\top})$
Phase II. $0.5<\mathsf{Tr}(\mathbf{I}_{r_{A}}-\mathbf{\Psi}_{t}\mathbf{\Psi}_{t}^{\top})<{r_{A}}-0.5$. Suppose that $\mathsf{Tr}\big((\mathbf{I}_{r_{A}}-\mathbf{\Lambda}_{t})\mathbf{\Lambda}_{t}\big)\geq\rho$, for some $\rho$ to be discussed shortly. Choosing $\eta\leq 0.5$, and $\eta={\cal O}\big(\frac{\rho(r-r_{A})^{2}}{r^{2}\kappa^{2}m}\big)$, it is straightforward to have
@@
Next, we show that $\rho\geq{\cal O}(\min\{\frac{1}{n},\frac{(r-r_{A})^{2}}{nr}\})\geq{\cal O}(\min\{\frac{1}{m},\frac{(r-r_{A})^{2}}{mr}\})$. Notice that $\mathsf{Tr}\big((\mathbf{I}_{r_{A}}-\tilde{\mathbf{\Lambda}}_{t})\tilde{\mathbf{\Lambda}}_{t}\big)\geq\sum_{i=1}^{r_{A}}\sigma_{i}^{2}(\mathbf{\Psi}_{t})(1-\sigma_{i}^{2}(\mathbf{\Psi}_{t}))\geq\sigma_{r_{A}}^{2}(\mathbf{\Psi}_{t})(1-\sigma_{r_{A}}^{2}(\mathbf{\Psi}_{t}))\geq{\cal O}(\min\{\frac{1}{n},\frac{(r-r_{A})^{2}}{nr}\})$, where the last inequality comes from the facts that i) for $x\in[a,b]$ with $0<a<0.5<b<1$, the smallest value of x(1x)x(1-x) is $\min\{a(1-a),b(1-b)\}$; and, ii) $\sigma_{r_{A}}^{2}(\mathbf{\Psi}_{t})$ belongs to interval $[a,b]$ with $a={\cal O}(\frac{(r-r_{A})^{2}}{nr})$ and $b=\frac{r_{A}-0.5}{r_{A}}=1-\frac{1}{2r_{A}}\leq 1-\frac{1}{n}$. Lemmas 15 and 11 are adopted to calculate $a$, that is, $a=\sigma_{r_{A}}^{2}(\mathbf{\Psi}_{t})\geq\sigma_{r_{A}}^{2}(\mathbf{\Psi}_{0})={\cal O}((r-r_{A})^{2}/nr)$.
@@ E.2.5 Convergence
Phase II. $0.5<\mathsf{Tr}(\mathbf{I}_{r_{B}}-\mathbf{\Phi}_{t}\mathbf{\Phi}_{t}^{\top})<{r_{B}}-0.5$. Suppose that $\mathsf{Tr}\big((\mathbf{I}_{r_{B}}-\mathbf{\Lambda}_{t}^{2})\mathbf{\Lambda}_{t}^{2}\big)\geq\rho$, for some $\rho$ to be discussed shortly. Choosing $\eta\leq 0.5$, and $\eta={\cal O}\big(\frac{\rho(r-r_{B})^{2}}{r^{2}\kappa^{2}m}\big)$, it is straightforward to have
@@
Next, we show that $\rho\geq{\cal O}(\min\{\frac{1}{m},\frac{(r-r_{B})^{2}}{mr}\})$. Notice that $\mathsf{Tr}\big((\mathbf{I}_{r_{B}}-\mathbf{\Lambda}_{t}^{2})\mathbf{\Lambda}_{t}^{2}\big)\geq\sum_{i=1}^{r_{B}}\sigma_{i}^{2}(\mathbf{\Phi}_{t})(1-\sigma_{i}^{2}(\mathbf{\Phi}_{t}))\geq\sigma_{r_{B}}^{2}(\mathbf{\Phi}_{t})(1-\sigma_{r_{B}}^{2}(\mathbf{\Phi}_{t}))\geq{\cal O}(\min\{\frac{1}{m},\frac{(r-r_{B})^{2}}{mr}\})$, where the last inequality comes from the facts that i) for $x\in[a,b]$ with $0<a<0.5<b<1$, the smallest value of x(1x)x(1-x) is $\min\{a(1-a),b(1-b)\}$; and, ii) $\sigma_{r_{B}}^{2}(\mathbf{\Phi}_{t})$ belongs to interval $[a,b]$ with $a={\cal O}\big(\frac{(r-r_{B})^{2}}{mr})$ and $b=\frac{r_{B}-0.5}{r_{B}}=1-\frac{1}{2r_{B}}\leq 1-\frac{1}{m}$. Lemmas 21 and 19 are adopted to calculate $a$, that is, $\sigma_{r_{B}}^{2}(\mathbf{\Phi}_{t})\geq\sigma_{r_{B}}^{2}(\mathbf{\Phi}_{0})={\cal O}((r-r_{B})^{2}/mr)$.